ย The hydrogen atom wave functions are written as ๐œ“๐‘›๐‘™๐‘š. State the values of n, l, and m. State the relation between a physical quantity and each quantum number. At t = 0 the hydrogen atom is in the superposition state ฮจ(๐‘Ÿโƒ—, 0) = 1 โˆš3 ๐œ“100 + ๐ด๐œ“210 where A is a real positive constant. Find A by normalization and determine the wave function at time ๐‘ก > 0. Find the average energy of the electron in eV given that the ground state energy is -13.6 eV.


1
Expert's answer
2020-08-04T16:00:38-0400

Given wave function is

"\psi(r,0) = \frac{1}{\sqrt {3}} \psi _{100}+ A \psi_{210}"

According to Normalization condition,

"\int |\psi(r,0)|^2 dr=1"

"\int |\frac{1}{\sqrt {3}} \psi _{100}+ A \psi_{210}|^2 dr=1"

"\frac{1}{3}\int |\psi _{100}|^2 dr + A^2 \int |\psi _{210}|^2 dr=1"

"\frac{1}{3}+A^2=1"

Since "\int |\psi _{100}|^2 dr =1 and" "\int |\psi _{210}|^2 dr=1"

On simplifying above equation, we get

"A=\sqrt\frac{2}{3}"

The wave function as a function of time is

"\psi(r,t) = \frac{1}{\sqrt {3}} \psi _{100} (r,t)+ \sqrt\frac{2}{3} \psi_{210}(r,t)"

"= \frac{1}{\sqrt {3}} \psi _{100} (r)e^{-\frac{iE_1t}{\hbar}}+ \sqrt\frac{2}{3} \psi_{210}(r)e^{-\frac{iE_2t}{\hbar}}"

The average energy of the electron is

"<E>=\frac{1}{3}<E_1>+\frac{2}{3}<E_2>"

"=\frac{1}{3}(-13.6 eV)+\frac{2}{3}(\frac{-13.6 eV}{2^2})"

"=-4.53 eV - 2.27 eV=-6.80 eV"